OpenStudy (anonymous):

5X+3Y=-11 7X-2Y=17 ORDER PAIR PLEASE?

myininaya (myininaya):

5X+3Y=-11 3Y=-11-5X (subtracted 5X on both sides) Y=-11/3-5X/3 (divided both sides by 3) the equation is 7X-2Y=17 (but we know what Y is in terms of X so we plug in -11/3-5X/3 for Y) 7X-2(-11/3-5X/3)=17 7X+22/3+10X/3=17 7X+10X/3=17-22/3 21X/3+10X/3=51/3-22/3 31X/3=29/3 X=29/3 * 3/31 X=29/31 but remember we solved that first equation for Y and it was in terms of just X so we can easily find what Y is by just plugging in what we found for X Since X=29/31 wherever we see a X in Y=-11/3-5X/3 we will put 29/31 so we have Y=-11/3-5(29/(31*3))

OpenStudy (anonymous):

ORDER PAIR?

OpenStudy (gw2011):

I assume that the solution you are looking for is in the format of: (29/31, -162/31)